After spending $1550 for tables and $3150 for chairs, a convention center manager finds that 40% of his original budget remains. find the amount that remains.
Answer by fractalier(2101)

If 40% of the budget remains, that means 60% was spent.
The 60% spent was the sum of $3150 and $1550, a total of $4700.
So we say 60 is to $4700 as 40 is to x and write it algebraically as
60/4700 = 40/x
We cross-multiply and solve
60x = 188000
x = 188000/60
x = $3133.33